Flying against the wind, a plane makes a 1050km trip in 3.5 hours. On the return trip, the flight takes 30 minutes less. What is the rate of the plane in calm air?
Answer by drk(1908)

This is an rtd problem with currents. Here is a table based on the given information:
plane . . . . . .rate . . . . . . . . . . . time . . . . .distance
against . . . .r-c=1050 . . . . . . . .3.5 . . . . . . 3675
with . . . . . .r+c . . . . . . . . .. . . .. 3 . . .. . . . .3675
we know
(i) r-c = 1050.
solve with for r+c, we get
(ii) r+c = 1225
adding (i) and (ii) together, we get
(iii) 2r = 2275
r = 1137.5
c then is 87.5 Kph